In a normal distribution 9.85% of the values are under 40 and 89.97% of the values are under 60. Find the point of inflection of the distribution
1
Expert's answer
2019-12-09T12:26:43-0500
"P(Z<Z_1)=0.0985"

"Z_1=\frac{40-\mu}{\sigma}=-1.29"

"P(Z<Z_2)=0.8997"

"Z_2=\frac{60-\mu}{\sigma}=1.28"

"\frac{60-\mu}{40-\mu}=-\frac{1.28}{1.29}"

"\mu=50.0"

Thus,


"\frac{60-50}{\sigma}=1.28"

"\sigma=7.8"

Points of inflection are


"x=50-7.8=42.2"

"x'=50+7.8=57.8"